# Privacy-aware Berrut Approximated Coded Computing for Federated Learning

### v1

@@ VI-A Computation with non-linear functions
We have compared our proposal (PBSS) to the original BSS without privacy (BSS) and the original BSS adding a differential privacy scheme (BSS + DP). As it was previously mentioned, we have tested our proposal with three non-linear functions that are often used as activation functions in machine learning solutions: ReLU (Fig. 3a), Sigmoid (Fig. 3b) and Swish (Fig. 4a). We have used the normalized relative mean error as performance metric
@@
Tables in Fig.3 show a summary of the the RME value when receiving different number of results from workers: $100$, $150$ and $200$, i.e. when the stragglers are $100$ nodes, $50$ nodes or 00 nodes, respectively (note that the number of workers $N=200$ is defined in Table I).
As expected, since ReLU and Swish are similar functions, their behaviours are also similar. In both cases PBSS and BSS improve the numerical precision when the number of received results increases. However, the results obtained when applying BSS plus differential privacy are not so good. They do not improve when the number of results is higher than $50$, offering less privacy ($\sigma=30$) with no guarantees against colluding nodes. Therefore, the original BSS proposal (without privacy) offers slightly better results than our proposal (PBSS), being this a little cost to pay to include privacy in the computation. Precisely, the cost can be measured in terms of percentage of accuracy loss, being approximately $0.008\%$ for ReLU, $0.071\%$ for Sigmoid, and $0.008\%$ for Swish.
@@ VI-B Testing Matrix Multiplications
Fig. 6 shows the precision results when using the common matrix multiplication scheme proposed in Section V-B using conventional matrices (Fig. 6a) and sparse matrices (Fig. 6b. In both cases, PBSS offers almost identical results than the original scheme without privacy (BSS), so adding privacy does not have a high cost in performance. Anyway, it is worthy to mention that, as expected, the lower the number of stragglers the better the result achieved. But, differently to the results of the previous analysis (non-linear functions and aggregation functions in Section VI-A, BSS and PBSS perform worse in presence of stragglers, so they needs from more results to reduce the error. This behaviour is even more accentuated with sparse matrices.
@@ VII Conclusions
In this paper we proposed a solution, coined as Private BACC Secret Sharring (PBSS), able to assure privacy in FL and, simultaneously, able to deal with three important aspects that other previous approaches in literature have not solved: (i) working properly with non-linear functions, (ii) working properly with large matrix multiplications, (iii) reduce the high communications overheads due to the presence of semi-hones nodes and (iv) manage stragglers nodes without a high impact on precision. PBSS operate with the model parameters in the coded domain, ensuring output privacy in the nodes without any additional communication cost. Besides, our proposal could be leveraged in secure federated learning for the training of a wide range of machine learning algorithm, since it does not depend on the specific algorithms and it is able to operate with large matrix multiplication.

### v2

@@ VI-A Computation with non-linear functions
We have compared our proposal (PBSS) to PBSS without privacy (BSS) and BSS adding a differential privacy scheme (BSS + DP). As it was previously mentioned, we have tested our proposal with three non-linear functions that are often used as activation functions in machine learning solutions: ReLU (Fig. 4a), Sigmoid (Fig. 4b) and Swish (Fig. 5a). We have used the normalized relative mean error as performance metric
@@
Tables in Fig.4 show a summary of the the RME value when receiving different numbers of results from workers: $100$, $150$ and $200$, i.e. when the stragglers are $100$ nodes, $50$ nodes or 00 nodes, respectively (note that the number of workers $N=200$ is defined in Table III).
As expected, since ReLU and Swish are similar functions, their behaviours are also similar. In both cases PBSS and BSS improve the numerical precision when the number of received results increases. However, the results obtained when applying BSS plus differential privacy are not so good. They do not improve when the number of results is higher than $50$, offering less privacy ($\sigma=30$) with no guarantees against colluding nodes. Therefore, the BSS scheme (without privacy) offers slightly better results than PBSS, being this a little cost to pay to include privacy in the computation. Precisely, the cost can be measured in terms of percentage of accuracy loss, being approximately $0.008\%$ for ReLU, $0.071\%$ for Sigmoid, and $0.008\%$ for Swish.
@@ VI-B Testing Matrix Multiplications
Fig. 7 shows the precision results when using the common matrix multiplication scheme proposed in Section V-B using conventional matrices (Fig. 7a) and sparse matrices (Fig. 7b. In both cases, PBSS offers almost identical results than the original scheme without privacy (BSS), so adding privacy does not have a high cost in performance. Anyway, it is worthy to mention that, as expected, the lower the number of stragglers the better the result achieved. But, differently to the results of the previous analysis (non-linear functions and aggregation functions in Section VI-A, BSS and PBSS perform worse in presence of stragglers, so they need more results to reduce the error. This behaviour is even more accentuated with sparse matrices.
@@ VII Conclusions
In this paper we proposed a solution, coined as Private BACC Secret Sharing (PBSS), able to assure privacy in FL and, simultaneously, able to deal with four important aspects that other previous approaches in literature have not solved: (i) working properly with non-linear functions, (ii) working properly with large matrix multiplications, (iii) reduce the high communications overheads due to the presence of semi-hones nodes and (iv) manage stragglers nodes without a high impact on precision. PBSS operate with the model parameters in the coded domain, ensuring output privacy in the nodes without any additional communication cost. Besides, our proposal could be leveraged in secure federated learning for the training of a wide range of machine learning algorithms, since it does not depend on the specific algorithms and it is able to operate with large matrix multiplications.